# Self-Foveate: Enhancing Diversity and Difficulty of Synthesized Instructions from Unsupervised Text via Multi-Level Foveation

### v1

@@ Appendix C Diversity Experiment
$\text{D}_{\text{EB}}$ $=1-\dfrac{1}{2\lvert\mathcal{X}_{\tau}\rvert}\sum_{x_{i}\in\mathcal{X}_{\tau}}\sum_{x_{j}\in\mathcal{X}_{\tau}}\dfrac{\phi(x_{i})\cdot\phi(x_{j})}{\lVert\phi(x_{i})\rVert^{2}\lVert\phi(x_{j})\rVert^{2}}$ (3)
@@
Appendix D Difficulty Head-to-Head Experiment
We prompt GPT-4o (gpt-4o-2024-08-06)(OpenAI, 2024) to determine which method-synthesized instruction set is more challenging. The prompts are shown below.
@@ Appendix D Difficulty Head-to-Head Experiment
Hyperparameters Values
frequency_penalty 0.0
max_completion_tokens 5
presence_penalty 0
temperature 0.5
top_p 1.0
@@
Appendix E Downstream Experiment
E.1 Software and Hardware Details
@@ E.2 Hyperparameters
Hyperparameters Values
cutoff_len 2048
learning_rate 0.0001
num_train_epochs 5.0
effective_batch_size 16
lr_scheduler_type cosine
max_grad_norm 1.0
warmup_steps 0
optim adamw_torch
quantization_bit 4
quantization_method bitsandbytes
lora_rank 8
lora_alpha 16
lora_dropout 0
@@
E.3 LLM Accuary Evaluation
LLM Accurary Evaluation System Prompt You are a fair judge. Your task is to determine if the generated answer correctly answers the question, even if it contains additional explanations. Rules: 1. The generated answer is correct if it contains the key information from the ground truth 2. Additional explanations or context in the generated answer should not make it incorrect 3. Only respond with ’Correct’ or ’Incorrect’
LLM Accurary Evaluation User Prompt Compare the following answers: Question: {question} Ground Truth Answer: {ground_truth} Generated Answer: {generated} Is the generated answer correct, regardless of any additional explanation? Respond only with ’Correct’ or ’Incorrect’.
Appendix F Answer Regeneration Evaluation
@@ F.2 Evaluation Hyperparameters
Hyperparameters Values
frequency_penalty 0.0
max_completion_tokens 50
presence_penalty 0
temperature 0.2
top_p 1.0
@@
Appendix G Use of AI Assistants
In this study, we utilized AI-powered tools, including ChatGPT and Grammarly, to enhance the linguistic accuracy of our manuscript through spell-checking and minor grammatical corrections. Additionally, the codebase was developed using Cursor to improve coding efficiency and accuracy.

### v2

@@ Appendix G Diversity Experiment
$\text{D}_{\text{EB}}$ $=1-\dfrac{2}{\lvert\mathcal{X}_{\tau}\rvert(\lvert\mathcal{X}_{\tau}\rvert-1)}\sum_{i<j}\dfrac{\phi(x_{i})\cdot\phi(x_{j})}{\lVert\phi(x_{i})\rVert\lVert\phi(x_{j})\rVert}$ (4)
@@
Appendix H Difficulty Head-to-Head Experiment
We prompt GPT-4o (gpt-4o-2024-08-06)(OpenAI et al., 2024) to determine which method-synthesized instruction set is more challenging. The prompts are shown below.
@@ Appendix H Difficulty Head-to-Head Experiment
Hyperparameters Values frequency_penalty 0.0 max_completion_tokens 5 presence_penalty 0 temperature 0.5 top_p 1.0
@@
Appendix I Downstream Experiment
I.1 Software and Hardware Details
@@ I.2 Hyperparameters
Hyperparameters Values cutoff_len 2048 learning_rate 0.0001 num_train_epochs 5.0 effective_batch_size 16 lr_scheduler_type cosine max_grad_norm 1.0 warmup_steps 0 optim adamw_torch quantization_bit 4 quantization_method bitsandbytes lora_rank 8 lora_alpha 16 lora_dropout 0
@@
I.3 LLM Accuracy Evaluation
LLM Accuracy Evaluation System Prompt You are a fair judge. Your task is to determine if the generated answer correctly answers the question, even if it contains additional explanations. Rules: 1. The generated answer is correct if it contains the key information from the ground truth 2. Additional explanations or context in the generated answer should not make it incorrect 3. Only respond with ’Correct’ or ’Incorrect’
LLM Accuracy Evaluation User Prompt Compare the following answers: Question: {question} Ground Truth Answer: {ground_truth} Generated Answer: {generated} Is the generated answer correct, regardless of any additional explanation? Respond only with ’Correct’ or ’Incorrect’.
Appendix J Answer Regeneration Evaluation
@@ I.3 LLM Accuracy Evaluation
Hyperparameters Values frequency_penalty 0.0 max_completion_tokens 50 presence_penalty 0 temperature 0.2 top_p 1.0
@@
Appendix K Use of AI Assistants
In this study, we utilized AI-powered tools, including ChatGPT, to improve the linguistic quality of the manuscript through spell-checking and minor grammatical corrections. Additionally, the codebase was developed using Cursor to enhance coding efficiency and accuracy.